What is the speed relative to the ground of a boat moving at 12 m/s relative to the river, which flows with 5 m/s right–angle to the boat's direction?



1
Expert's answer
2020-12-21T13:53:20-0500

"\overrightarrow{v}_a=\overrightarrow{v}_r+\overrightarrow{v}_e"


"v_a=\sqrt{v_r^2+v_e^2}=\sqrt{12^2+5^2}=13(m/s)"